What is the equation of a parabola with vertex (0,0) and focus (0,4)?
Answer by stanbon(75887)

What is the equation of a parabola with vertex (0,0) and focus (0,4)?
---------
You might derive the equation or follow a form.
-------------
Following a formula you have h = 0, k = 0, p = 4
Form: (x-h)^2 = 4p(y-k)
Substitute the values:
(x-0)^2 = 4*4(y-0)
16y = x^2
y = (1/16)x^2
==============
Cheers,
Stan H.